# A Host–Kra $\mathbb{F}_{2}^{\omega}$-system of order $5$ that is not Abramov of order $5$, and non-measurability of the inverse theorem for the $U^{6}(\mathbb{F}_{2}^{n})$ norm

### v2

@@ 4. Two key nilspaces
Now we verify ergodicity. Let $(Q,S)\colon\mathbb{F}_{2}\to X_{5,r}$ be an arbitrary map. Then $Q$ is linear, so certainly $d^{3}Q=0$. Since $r\leq 5$, every map $S\colon\mathbb{F}_{2}\to\frac{1}{2^{r}}\mathbb{Z}/\mathbb{Z}$ lies in $\mathrm{Poly}^{6}(\mathbb{F}_{2}\to\frac{1}{2^{r}}\mathbb{Z}/\mathbb{Z})$ by Lemma A.23, and hence by Lemma 4.1 all pairs $(Q,S)$ lie in $C^{1}(X_{5,r})$, giving the claim.
Now we verify the corner completion axiom. Let $(Q,S)\colon\mathbb{F}_{2}^{n}\backslash\{1\}^{n}\to X_{5,r}$ be a map such that the restriction of $(Q,S)$ to any $(n-1)$-face of $\{0,1\}^{n}\equiv\mathbb{F}_{2}^{n}$ containing $0^{n}$ is in $C^{n-1}(X_{5,r})$. From the corner completion property of $X_{2}$, we may extend $Q$ to an $n$-cube $Q\colon\mathbb{F}_{2}^{n}\to X_{2}$, and then by Lemma 4.1 we can find a lift $(Q,S^{\prime})\colon\mathbb{F}_{2}^{n}\to X_{5,r}$ which is an $n$-cube. By (1), we conclude that the difference $S-S^{\prime}\colon\mathbb{F}_{2}^{n}\backslash\{1\}^{n}\to\frac{1}{2^{r}}\mathbb{Z}/\mathbb{Z}$ is a degree $5$ polynomial on each $(n-1)$-face of $\{0,1\}^{n}\equiv\mathbb{F}_{2}^{n}$ containing $0^{n}$. By the corner completion property of ${\mathcal{D}}^{5}(\frac{1}{2^{r}}\mathbb{Z}/\mathbb{Z})$, we may extend $S-S^{\prime}$ to a degree $5$ polynomial from $\mathbb{F}_{2}^{n}$ to $\frac{1}{2^{r}}\mathbb{Z}/\mathbb{Z}$; the resulting extension $S\colon\mathbb{F}_{2}^{n}\to\frac{1}{2^{r}}\mathbb{Z}/\mathbb{Z}$ then obeys (1), so that $(Q,S)$ is now extended to an $n$-cube on $X_{5,r}$ as required. When $n=6$, the extension of $Q$ is unique, and equation (1) (with $x=0$ and $h_{1},\dots,h_{6}$ the standard basis) also shows that the extension of $S$ is unique, so that $X_{5,r}$ is $5$-step as claimed.
@@
for all $6$-cubes $(q_{\omega},t_{\omega})_{\omega\in\{0,1\}^{6}}$ in $X_{2}\times\frac{1}{2^{r-1}}\mathbb{Z}/\mathbb{Z}$ (one can check that $02\tilde{\rho}=$ 0, so that this cocycle does indeed take values in $\frac{1}{2}\mathbb{Z}/\mathbb{Z}$). As we will not need this description of $X_{5,r}$ here, we leave the justification of this claim to the interested reader.
@@
This gives a negative answer (in the case $p=2,k=5$) to [4, Question 5.18], which asked the more general question of whether every $k$-step compact $p$-homogeneous nilspace has an injective nilspace morphism into a $k$-step compact filtered abelian group. As noted in that paper, an affirmative answer to this question for a given value of $p$ and $k$ would imply an affirmative answer to Conjecture 1.2 (and hence Conjecture 1.3 and Conjecture 1.1) for those values of $p,k$. Indeed, [4, Question 5.18] was answered affirmatively for $k\leq p+1$, leading to the corresponding results on Conjectures 1.2, 1.3, 1.1 mentioned in the introduction. Thus, Proposition 4.5 can be viewed as a weaker version of Theorem 1.6.
@@
of the polynomial $P$. We claim that this derivative is constant, by the following standard argument. If $(q_{0},s_{0}),(q_{1},s_{1})\in X_{5,1}$, then the tuple $(Q,S)\colon\mathbb{F}_{2}^{6}\to X_{5,1}$ defined by
$(Q,S)(\omega)\coloneqq\left(q_{\omega_{1}},s_{\omega_{1}}+1_{\omega_{2}=\dots=\omega_{k+1}=0}\frac{1}{2}\right)$
can easily be verified to obey the axioms (i), (ii) required to be a $6$-cube in $X_{5,1}$. From the quintic nature of $P$ we conclude that
@@
On the other hand, the function $(q,s)\mapsto e(s)$ is also an eigenfunction of this operator with eigenvalue $e(\frac{1}{2})$. Since the Koopman operator $V^{u}$ is unitary, and $e(P)$ has a non-zero inner product with $e(s)$, the eigenvalue of $e(P)$ must also be $e(\frac{1}{2})$, thus
@@
While the above proposition shows that $X_{5,1}$ cannot be embedded into a finite filtered abelian group, [4, Theorem 1.7] does show that there is a fibration $\pi\colon Y\to X_{5,1}$ (as defined in [16, Definition 7.1], [3, Definition 3.3.7]) that has the structure of a finite filtered abelian group and has good lifting properties; this result was in particular used in [4] to give an alternate proof of Conjecture 1.1 in both high and low characteristic. In fact, we can explicitly give such an extension. Let $G$ denote the abelian group $\mathbb{Z}/4\mathbb{Z}$ with the degree $3$ filtration
@@ 5.1. Constructing the counterexample
for some absolute constant $\eta>0$ and some Fourier coefficients $c_{\omega}\in X_{2}$ (which may depend on $n$ and $S$), using the usual $\mathbb{F}_{2}$-valued inner product
@@
on the vector space $X_{2}$. Applying the Cauchy–Schwarz–Gowers inequality (see e.g., [11, (5.5)]) we conclude that
@@ 5.2. Equidistribution theory for $Q,S$
for all $a_{1},\dots,a_{M},b_{1},\dots,b_{d}\in\mathbb{F}_{2}$. This is the composition of the nilspace morphism $(Q,S)\colon{\mathcal{D}}^{1}(\mathbb{F}_{2}^{n})\to X_{5,5}$ with a (random) linear transformation from $\mathbb{F}_{2}^{M+d}$ to $\mathbb{F}_{2}^{n}$, and so $(\tilde{Q},\tilde{S})$ is a (random) nilspace morphism from ${\mathcal{D}}^{1}(\mathbb{F}_{2}^{M+d})$ to $X_{5,5}$, or equivalently a (random) $M+d$-cube in $X_{5,5}$. Also, regardless of the choice of sampling vectors $v_{1},\dots,v_{d}$, $(\tilde{Q},\tilde{S})$ must agree with $(Q,S)$ on $\mathbb{F}_{2}^{M}$ in the sense that
@@
denote the space of all $M+d$-cubes $(\tilde{Q},\tilde{S})$ that agree with the $M$-cube $(Q_{0},S_{0})$ on the face $\mathbb{F}_{2}^{M}\times\{0^{d}\}$ in the sense of (32); this is a non-empty finite set whose cardinality is bounded uniformly in $n$. For each choice of $(Q,S),d$, let $\mu_{Q,S}^{(d)}$ denote the distribution of the random variable $(\tilde{Q},\tilde{S})$ generated by the random variables $v_{1},\dots,v_{d}$, thus $\mu_{Q,S}^{(d)}$ is the probability measure on $\Sigma^{(d)}_{Q_{0},S_{0}}$ defined by the formula
@@
with probability $1-o(1)$, where $o(1)$ denotes any quantity that goes to zero as $n\to\infty$ holding all other parameters not depending on $n$ (such as $d$) fixed. Here $d_{\operatorname{TV}}$ denotes the total variation distance between probability measures.
Informally, this theorem asserts that the condition (32) on the $M+d$-cube $(\tilde{Q},\tilde{S})$ is asymptotically the only constraint that could control (or even bias) the distribution of this $M+d$-cube. One could replace the total variation distance here by any other reasonable metric, since $\mu_{Q,S}^{(d)}$, $\overline{\mu}_{Q_{0},S_{0}}^{(d)}$ are supported on finite sets of cardinality bounded uniformly on $n$.
@@
In particular, all fibers have the same cardinality, and hence the uniform measure $\overline{\mu}_{Q_{0},S_{0}}^{(d)}$ pushes forward to the uniform measure $\overline{\mu}_{Q_{0}}^{(d)}$. Also, by definition the sampling measure $\mu_{Q,S}^{(d)}$ pushes forward to the sampling measure $\mu_{Q}^{(d)}$. Hence, in order to establish (34) with probability $1-o(1)$, a natural first step would be to first show the weaker claim that
@@
By the finite Fourier transform, it thus suffices to establish the claim
[…]
with probability $1-o(1)$ for any fixed non-trivial character $\xi\colon H\to\mathbb{T}$, where $\tilde{Q}_{*}=\tilde{Q}_{*}(Q)$ is an arbitrary element of $\Sigma^{(d)}_{Q}$ (the exact choice is unimportant as it does not affect the magnitude of the left-hand side). By Chebyshev’s inequality, it suffices to show that
[…]
@@
Since $d$ is fixed and $n$ is going to infinity, we see that the vectors $v_{1},\dots,v_{d}$, $v^{\prime}_{1},\dots,v^{\prime}_{d}$, $e_{1},\dots,e_{M}$ will be linearly independent with probability $1-o(1)$. Hence we may restrict to this portion of the average with acceptable error. Applying a linear change of variables (which does not affect the distribution of the random variable $Q$), we may then normalize $v_{i}=e_{M+i}$ and $v^{\prime}_{i}=e_{M+d+i}$ for $i=1,\dots,d$. It will thus suffice to show that
@@
for all quadratic polynomials $Q\colon\mathbb{F}_{2}^{n}\to X_{2}$. But if we let $P\colon\mathbb{F}_{2}^{M+d}\to X_{2}$ be an element of the group $H$ that is not annihilated by $\xi$, one easily checks that the function $Q\colon\mathbb{F}_{2}^{n}\to X_{2}$ defined by
@@
is a quadratic polynomial for which the left-hand side of (37) is non-zero. Thus we have the desired equidistribution (36).
@@
with probability $1-o(1)$, where $1_{\tilde{Q}=\tilde{Q}_{*}}$ denotes the indicator function to the set $\{(\tilde{Q},\tilde{S})\in\Sigma^{(d)}_{Q_{0},S_{0}}:\tilde{Q}=\tilde{Q}_{*}\}$. Note from (36) that with probability $1-o(1)$, both of these measures differ in mass by $o(1)$. Once one fixes $\tilde{Q}=\tilde{Q}_{*}$, the variable $\tilde{S}$ ranges in a coset $\tilde{S}_{(S_{0},\tilde{Q}_{*})}+K$ of the finite abelian group $K$ defined in (35), where we arbitrarily choose one representative $\tilde{S}_{(S_{0},\tilde{Q}_{*})}$ of this coset for each choice of $S_{0},\tilde{Q}_{*}$. By Fourier analysis, it thus suffices to show that
[…]
with probability $1-o(1)$ for each non-trivial character $\xi\colon K\to\mathbb{T}$ of $K$. As before, it suffices by the Chebyshev inequality to show that
[…]
@@
Suppose that Conjecture 1.3 holds for $p=2$ and $k=5$, and let $\varepsilon\colon\mathbb{N}\to\mathbb{R}^{+}$ be a function decreasing to zero. Let $M$ be as in Proposition 5.3. Then there exists a sequence of $n$ going to infinity, an integer $1\leq m\leq M$ and a function $\Phi\colon(\frac{1}{2^{5}}\mathbb{Z}/\mathbb{Z})^{\mathbb{F}_{2}^{M}}\to\mathbb{T}$ such that, for $n$ along some sequence going to infinity, and for each $n$ in this sequence we may find a deterministic $n$-cube $(Q,S)$ in $X_{5,5}$, with the resulting $M$-cube $(Q_{0},S_{0})$ (and hence $\Sigma^{(d)}_{Q_{0},S_{0}}$ and $\overline{\mu}_{Q_{0},S_{0}}^{(d)}$) independent of $n$, obeying the estimates
(39) $|\mathbb{E}_{x\in\mathbb{F}_{2}^{n}}e(P(x))-e(\Phi((S(x+(a,0^{n-M})))_{a\in\mathbb{F}_{2}^{M}}))|\leq 2\varepsilon(m)$
@@
(40) $|\mathbb{E}_{x\in\mathbb{F}_{2}^{n}}e(S(x)-P(x))|\geq\frac{1}{m},$
such that $\mu_{Q,S}^{(d)}$ converges in total variation norm to $\overline{\mu}_{Q_{0},S_{0}}^{(d)}$ for each $d\geq 0$.
@@
Applying Theorem 5.4 and a standard diagonal argument, we obtain along a sequence $n$ going to infinity, an $n$-cube $(Q,S)$ in $X_{5,5}$, an integer $1\leq m\leq M$, a polynomial $P\in\mathrm{Poly}^{5}(\mathbb{F}_{2}^{n})$, and a function $\Phi\colon(\frac{1}{2^{5}}\mathbb{Z}/\mathbb{Z})^{\mathbb{F}_{2}^{M}}\to\mathbb{T}$ obeying (30), (31) such that
@@
By either Remark A.3 or A.4, this space has the structure of a finite $5$-step $2$-homogeneous nilspace (it is easy to see that the two nilspace structures given by these remarks agree). This space will not be ergodic in general, so the equivalence relation $\sim_{0}$ on this space introduced in Remark A.2 can be non-trivial. The morphism $(Q_{0},S_{0})$ is a point in $\operatorname{Hom}_{\Box}(\mathbb{F}_{2}^{M},X_{5,5})$, and we define $X_{(Q_{0},S_{0})}$ to be the equivalence class of this point:
@@
This is then an ergodic finite $5$-step $2$-homogeneous nilspace.
@@
for all $(\tilde{Q},\tilde{S})\in\Sigma^{(1+s)}_{Q_{0},S_{0}}$; thus $\pi_{s}(\tilde{Q},\tilde{S})$ is the tuple formed by restricting $(\tilde{Q},\tilde{S})$ to the affine subspaces $\mathbb{F}_{2}^{M}\times(1,\omega)$ of $\mathbb{F}_{2}^{M+1+s}$ for $\omega\in\{0,1\}^{s}$. Let us first check that $\pi_{s}(\tilde{Q},\tilde{S})$ lies in $C^{s}(X_{(Q_{0},S_{0})})$ as claimed. Since $(\tilde{Q},\tilde{S})$ is a $M+1+s$-cube in $X_{5,5}$, the map
@@
is an $s$-cube in $C^{M}(X_{5,5})$. Applying (62), the tuple $\pi_{s}(\tilde{Q},\tilde{S})$ is thus a $s$-cube in $C^{M}(X_{5,5})$, and thus lies in a single equivalence class of $\sim_{0}$. A similar argument shows that the pair
@@
is a $1$-cube in $C^{M}(X_{5,5})$, and so the two elements of this pair are also equivalent by $\sim_{0}$. By (32), the first map is $(Q_{0},S_{0})$, and hence $\psi_{s}(\tilde{Q},\tilde{S})$ is an $s$-cube in $X_{(Q_{0},S_{0})}$ as claimed.
Next, we claim that the map $\pi_{s}$ is surjective. Let $((Q^{\prime}_{\omega},S^{\prime}_{\omega}))_{\omega\in\{0,1\}^{s}}$ be an $s$-cube in $X_{(Q_{0},S_{0})}$. Our goal is to locate an $M+1+s$-cube $(\tilde{Q},\tilde{S})$ in $X_{5,5}$ such that
@@
for all $a\in\mathbb{F}_{2}^{M}$ and $\omega\in\{0,1\}^{s}$. So $(\tilde{Q},\tilde{S})$ is already partially specified on the set
@@
Now we claim that all the fibers of $\pi_{s}$ have the same cardinality. Observe that if $(\tilde{Q},\tilde{S}),(\tilde{Q}^{\prime},\tilde{S}^{\prime})\in\Sigma^{(1+s)}_{Q_{0},S_{0}}$ have the same image under $\pi_{s}$, then $\tilde{Q}-\tilde{Q}^{\prime}$ is an element of $\mathrm{Poly}^{2}(\mathbb{F}_{2}^{M+1+s}\to X_{2})$ that vanishes on the set (41); and if $\tilde{Q}=\tilde{Q}^{\prime}$, then $\tilde{S}-\tilde{S}^{\prime}$ is an element of $\mathrm{Poly}^{5}(\mathbb{F}_{2}^{M+1+s}\to\frac{1}{2^{5}}\mathbb{Z}/\mathbb{Z})$ that vanishes on (41). Conversely, if $(\tilde{Q},\tilde{S})\in\Sigma^{(1+s)}_{Q_{0},S_{0}}$ and $\tilde{S}-\tilde{S}^{\prime}$ is an element of $\mathrm{Poly}^{5}(\mathbb{F}_{2}^{M+1+s}\to\frac{1}{2^{5}}\mathbb{Z}/\mathbb{Z})$ that vanishes on (41), then $(\tilde{Q},\tilde{S}^{\prime})$ is an element of $\Sigma^{(1+s)}_{Q_{0},S_{0}}$ with the same image as $(\tilde{Q},\tilde{S})$ under $\pi_{s}$. To conclude the claim, it suffices to show that whenever $(\tilde{Q},\tilde{S})\in\Sigma^{(1+s)}_{Q_{0},S_{0}}$ and $\tilde{Q}-\tilde{Q}^{\prime}$ is an element of $\mathrm{Poly}^{2}(\mathbb{F}_{2}^{M+1+s}\to X_{2})$ that vanishes on (41), then there exists $(\tilde{Q}^{\prime},\tilde{S}^{\prime})\in\Sigma^{(1+s)}_{Q_{0},S_{0}}$ with the same image as $(\tilde{Q},\tilde{S})$ under $\pi_{s}$. By Lemma 4.1, we can at least find a function $\tilde{S}^{\prime\prime}\colon\mathbb{F}_{2}^{n}\to\frac{1}{2}\mathbb{Z}/\mathbb{Z}$ with $(\tilde{Q}^{\prime},\tilde{S}^{\prime\prime})$ an $n$-cube in $X_{5,5}$. If the $\tilde{S}^{\prime\prime}-\tilde{S}$ vanished on (41), we would be done; but the best that can be said at present is that this function is a polynomial of degree $k$ on (42) and on (43), again thanks to Lemma 4.1. Applying the completion axiom (or [3, Lemma 3.1.5]) many times, we can then find $P\in\mathrm{Poly}^{5}(\mathbb{F}_{2}^{n}\to\frac{1}{2^{5}}\mathbb{Z}/\mathbb{Z})$ which agrees with $\tilde{S}^{\prime\prime}-\tilde{S}$ on (41); setting $\tilde{S}^{\prime}\coloneqq\tilde{S}^{\prime\prime}-P$ gives the claim.
@@
Suppose that Conjecture 1.3 holds for $p=2$ and $k=5$, and let $\varepsilon\colon\mathbb{N}\to\mathbb{R}^{+}$ be a function decreasing to zero. Let $M$ be as in Proposition 5.3, and let $n,(Q,S),(Q_{0},S_{0}),m,\Phi$ be as in Corollary 5.5. If for any $s\geq 0$ we select $x,h_{1},\dots,h_{s}\in\mathbb{F}_{2}^{n}$ uniformly and independently at random, then the random tuple
$((a\mapsto(Q,S)((a,0^{n-M})+x+\sum_{i=1}^{s}\omega_{i}h_{i})))_{\omega\in\{0,1\}^{s}}$

### v3

@@ 4. Two key nilspaces
Now we verify ergodicity. Let $(Q,S)\colon\mathbb{F}_{2}\to X_{5,r}$ be an arbitrary map. Then $Q$ is linear, so certainly $d^{3}Q=0$. Since $r\leq 5$, every map $S\colon\mathbb{F}_{2}\to\frac{1}{2^{r}}\mathbb{Z}/\mathbb{Z}$ lies in $\mathrm{Poly}^{5}(\mathbb{F}_{2}\to\frac{1}{2^{r}}\mathbb{Z}/\mathbb{Z})$ by Lemma A.23, and hence by Lemma 4.1 all pairs $(Q,S)$ lie in $C^{1}(X_{5,r})$, giving the claim.
[…]
@@
for all $6$-cubes $(q_{\omega},t_{\omega})_{\omega\in\{0,1\}^{6}}$ in $X_{2}\times\frac{1}{2^{r-1}}\mathbb{Z}/\mathbb{Z}$ (one can check that $2\tilde{\rho}=0$, so that this cocycle does indeed take values in $\frac{1}{2}\mathbb{Z}/\mathbb{Z}$). As we will not need this description of $X_{5,r}$ here, we leave the justification of this claim to the interested reader.
@@
This gives a negative answer (in the case $p=2,k=5$) to [5, Question 5.18], which asked the more general question of whether every $k$-step compact $p$-homogeneous nilspace has an injective nilspace morphism into a $k$-step compact filtered abelian group. As noted in that paper, an affirmative answer to this question for a given value of $p$ and $k$ would imply an affirmative answer to Conjecture 1.2 (and hence Conjecture 1.3 and Conjecture 1.1) for those values of $p,k$. Indeed, [5, Question 5.18] was answered affirmatively for $k\leq p+1$, leading to the corresponding results on Conjectures 1.2, 1.3, 1.1 mentioned in the introduction. Thus, Proposition 4.5 can be viewed as a weaker version of Theorem 1.6.
@@
of the polynomial $P$. We claim that this derivative is constant. If $(q_{0},s_{0}),(q_{1},s_{1})\in X_{5,1}$, then the tuple $(Q,S)\colon\mathbb{F}_{2}^{6}\to X_{5,1}$ defined by
$(Q,S)(\omega)\coloneqq\left(q_{\omega_{1}},s_{\omega_{1}}+1_{\omega_{2}=\dots=\omega_{6}=0}\frac{1}{2}\right)$
can be verified to obey the axioms (i), (ii) required to be a $6$-cube in $X_{5,1}$ by checking the conditions in the description (27). From the quintic nature of $P$ we conclude that
@@
On the other hand, the function $(q,s)\mapsto e(s)$ is also an eigenfunction of this operator with eigenvalue $e(\frac{1}{2})$. Since the Koopman operator $V^{u}$ is unitary, therefore its eigenvectors with distinct eigenvalues are orthogonal, and $e(P)$ has a non-zero inner product with $e(s)$, the eigenvalue of $e(P)$ must also be $e(\frac{1}{2})$, thus
@@
While the above proposition shows that $X_{5,1}$ cannot be embedded into a finite filtered abelian group, [5, Theorem 1.7] does show that there is a fibration $\pi\colon Y\to X_{5,1}$ (as defined in [17, Definition 7.1], [3, Definition 3.3.7]) such that $Y$ has the structure of a finite filtered abelian group and $\pi$ has good lifting properties; this result was in particular used in [5] to give an alternate proof of Conjecture 1.1 in both high and low characteristic. In fact, we can explicitly give such an extension. Let $G$ denote the abelian group $\mathbb{Z}/4\mathbb{Z}$ with the degree $3$ filtration
@@ 5.1. Constructing the counterexample
for some absolute constant $\eta>0$ and some coordinates of Fourier frequencies $c_{\omega}\in X_{2}$ (which may depend on $n$ and $S$), using the usual $\mathbb{F}_{2}$-valued inner product
@@
on the vector space $X_{2}$. Applying the Cauchy–Schwarz–Gowers inequality (see e.g., [12, (5.5)]) we conclude that
@@ 5.2. Equidistribution theory for $Q,S$
[…]
@@
$\int_{\Sigma^{(d)}_{Q_{0},S_{0}}}G(\tilde{Q},\tilde{S})\ d\mu_{Q,S}^{(d)}(\tilde{Q},\tilde{S})=\mathbb{E}_{v_{1},\dots,v_{d}\in\mathbb{F}_{2}^{n}}G(\tilde{Q}_{Q,v_{1},\dots,v_{d}},\tilde{S}_{S,v_{1},\dots,v_{d}})$
@@
with probability 1o(1)1-o(1), where $o(1)$ denotes any quantity that goes to zero as $n\to\infty$ holding all other parameters not depending on $n$ (such as $d$) fixed. Here $d_{\operatorname{TV}}$ denotes the total variation distance between probability measures.
[…]
@@
In particular, all fibers have the same cardinality, and hence the uniform measure $\overline{\mu}_{Q_{0},S_{0}}^{(d)}$ pushes forward to the uniform measure $\overline{\mu}_{Q_{0}}^{(d)}$. Also, by definition the sampling measure $\mu_{Q,S}^{(d)}$ pushes forward to the sampling measure $\mu_{Q}^{(d)}$. Hence, in order to establish (37) with probability 1o(1)1-o(1), a natural first step would be to first show the weaker claim that
@@
By the Diaconis–Shahshahani Upper Bound Lemma [^8] For a probability measure $\mu$ on a finite abelian group $H$ (or any coset thereof), the Diaconis–Shahshahani Upper Bound Lemma is the inequality $d_{\operatorname{TV}}(\mu,u)^{2}\leq\tfrac{1}{4}\sum_{\xi\neq 0}|\widehat{\mu}(\xi)|^{2},$ where $u$ denotes the uniform measure and $\widehat{\mu}(\xi)$ are the Fourier coefficients with respect to the characters of $H$. (cf. [7, Lemma 1]), it thus suffices to establish the claim
$\int_{\Sigma^{(d)}_{Q_{0}}}e(\xi\cdot(\tilde{Q}-\tilde{Q}_{*}))\ d\mu_{Q}^{(d)}=o(1)$
with probability 1o(1)1-o(1) for any fixed non-trivial character $\xi\colon H\to\mathbb{T}$, where $\tilde{Q}_{*}=\tilde{Q}_{*}(Q)$ is an arbitrary element of $\Sigma^{(d)}_{Q}$ (the exact choice is unimportant as it does not affect the magnitude of the left-hand side). By Chebyshev’s inequality, it suffices to show that
$\mathbb{E}_{Q,S}\left|\int_{\Sigma^{(d)}_{Q_{0}}}e(\xi\cdot(\tilde{Q}-\tilde{Q}_{*}))\ d\mu_{Q}^{(d)}\right|^{2}=o(1).$
@@
Since $d$ is fixed and $n$ is going to infinity, we see that the vectors $v_{1},\dots,v_{d}$, $v^{\prime}_{1},\dots,v^{\prime}_{d}$, $e_{1},\dots,e_{M}$ will be linearly independent with probability 1o(1)1-o(1). Hence we may restrict to this portion of the average with acceptable error. Applying a linear change of variables (which does not affect the distribution of the random variable $Q$), we may then normalize $v_{i}=e_{M+i}$ and $v^{\prime}_{i}=e_{M+d+i}$ for $i=1,\dots,d$. It will thus suffice to show that
@@
for all quadratic polynomials $Q\colon\mathbb{F}_{2}^{n}\to X_{2}$. But if we let $P\colon\mathbb{F}_{2}^{M+d}\to X_{2}$ be an element of the group $H$ that is not annihilated by $\xi$, then the function $Q\colon\mathbb{F}_{2}^{n}\to X_{2}$ defined by
@@
is a quadratic polynomial [^9] The composition of a homomorphism, here the projection from $\mathbb{F}_{2}^{n}$ onto $\mathbb{F}_{2}^{M+d}$, with a polynomial of degree $d$ is again a polynomial of degree $d$. for which the left-hand side of (40) is non-zero. Thus we have the desired equidistribution (39).
@@
with probability 1o(1)1-o(1), where $1_{\tilde{Q}=\tilde{Q}_{*}}$ denotes the indicator function to the set $\{(\tilde{Q},\tilde{S})\in\Sigma^{(d)}_{Q_{0},S_{0}}:\tilde{Q}=\tilde{Q}_{*}\}$. Note from (39) that with probability 1o(1)1-o(1), both of these measures differ in mass by $o(1)$. Once one fixes $\tilde{Q}=\tilde{Q}_{*}$, the variable $\tilde{S}$ ranges in a coset $\tilde{S}_{(S_{0},\tilde{Q}_{*})}+K$ of the finite abelian group $K$ defined in (38), where we arbitrarily choose one representative $\tilde{S}_{(S_{0},\tilde{Q}_{*})}$ of this coset for each choice of $S_{0},\tilde{Q}_{*}$. By Fourier analysis, it thus suffices to show that
$\int_{\Sigma^{(d)}_{Q_{0},S_{0}}}e(\xi\cdot(\tilde{S}-\tilde{S}_{(S_{0},\tilde{Q}_{*})}))1_{\tilde{Q}=\tilde{Q}_{*}}\ d\mu_{Q,S}^{(d)}(\tilde{Q},\tilde{S})=o(1)$
with probability 1o(1)1-o(1) for each non-trivial character $\xi\colon K\to\mathbb{T}$ of $K$. As before, it suffices by the Chebyshev inequality to show that
$\mathbb{E}_{Q,S}|\int_{\Sigma^{(d)}_{Q,S}}e(\xi\cdot(\tilde{S}-\tilde{S}_{S_{0},\tilde{Q}_{*}}))1_{\tilde{Q}=\tilde{Q}_{*}}\ d\mu_{Q,S}^{(d)}(\tilde{Q},\tilde{S})|^{2}=o(1).$
@@
Suppose that Conjecture 1.3 holds for $p=2$ and $k=5$, and let $\varepsilon\colon\mathbb{N}\to\mathbb{R}^{+}$ be a function decreasing to zero. Let $M$ be as in Proposition 5.3. Then there exist an integer $1\leq m\leq M$, a function $\Phi\colon(\frac{1}{2^{5}}\mathbb{Z}/\mathbb{Z})^{\mathbb{F}_{2}^{M}}\to\mathbb{T}$, an $M$-cube $(Q_{0},S_{0})$ (which also defines $\Sigma^{(d)}_{Q_{0},S_{0}}$ and $\overline{\mu}_{Q_{0},S_{0}}^{(d)}$), and an infinite sequence of dimensions $n\to\infty$ such that the following holds. For each $n$ in this sequence, there are a deterministic $n$-cube $(Q,S)$ in $X_{5,5}$, with associated $M$-cube $(Q_{0},S_{0})$, such that one has
(42) $\Big{|}\mathbb{E}_{x\in\mathbb{F}_{2}^{n}}e(P(x))-e\Big{(}\Phi\big{(}(S(x+(a,0^{n-M})))_{a\in\mathbb{F}_{2}^{M}}\big{)}\Big{)}\Big{|}\leq 2\varepsilon(m),$
@@
(43) $\Big{|}\mathbb{E}_{x\in\mathbb{F}_{2}^{n}}e(S(x)-P(x))\Big{|}\geq\frac{1}{m},$
with $\mu_{Q,S}^{(d)}$ converging in total variation norm to $\overline{\mu}_{Q_{0},S_{0}}^{(d)}$ for each $d\geq 0$.
@@
Applying Proposition 5.3, Theorem 5.4, and a standard diagonal argument, we obtain along a sequence $n$ going to infinity, an $n$-cube $(Q,S)$ in $X_{5,5}$, an integer $1\leq m\leq M$, a polynomial $P\in\mathrm{Poly}^{5}(\mathbb{F}_{2}^{n})$, and a function $\Phi\colon(\frac{1}{2^{5}}\mathbb{Z}/\mathbb{Z})^{\mathbb{F}_{2}^{M}}\to\mathbb{T}$ obeying (33), (34) such that
@@
By either Remark A.3 or A.4, this space has the structure of a finite $5$-step 22-homogeneous nilspace (it is easy to see that the two nilspace structures given by these remarks agree). This space will not be ergodic in general, so the equivalence relation $\sim_{0}$ on this space introduced in Remark A.2 can be non-trivial. The morphism $(Q_{0},S_{0})$ is a point in $\operatorname{Hom}_{\Box}(\mathbb{F}_{2}^{M},X_{5,5})$, and we define $X_{(Q_{0},S_{0})}$ to be the equivalence class of this point:
@@
This is then an ergodic finite $5$-step 22-homogeneous nilspace.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Suppose that Conjecture 1.3 holds for $p=2$ and $k=5$, and let $\varepsilon\colon\mathbb{N}\to\mathbb{R}^{+}$ be a function decreasing to zero. Let $M$ be as in Proposition 5.3, and let $n,(Q,S),(Q_{0},S_{0}),m,\Phi$ be as in Corollary $5.$ If for any $s\geq 0$ we select $x,h_{1},\dots,h_{s}\in\mathbb{F}_{2}^{n}$ uniformly and independently at random, then the random tuple
$\bigl{(}(a\mapsto(Q,S)((a,0^{n-M})+x+\sum_{i=1}^{s}\omega_{i}h_{i}))\bigr{)}_{\omega\in\{0,1\}^{s}}$